OpenStudy (anonymous):

Find the point on the curve y = x^2 closest to the point (0, 1).

OpenStudy (anonymous):

Calculus?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

There are two answers, due to the symmetry of the function. For now, we can assume x positive.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Find the distance from (0,1) to the curve, and minimize that function.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We need to set the distance up as a function of x. This is a job for the Pythagorean theorem.\[(d(x))^2=(x-x_1)^2+(y-y_1)^2=x^2+(x^2-1)^2\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

When you multiply the second term, you will get an equation that is quadratic in form (fourth degree polynomial). Take the derivative, set it to zero, and solve for x. Once you have x, you can get y.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Note that we worked with the square of the distance. If we minimize it, the distance will also be minimized.

OpenStudy (anonymous):

ok got you

OpenStudy (anonymous):

\[Let f(x)=(d(x))^2 \implies f(x)=x^4-x^2+1 \implies f'(x)=4x^3-2x=0\]\[\implies 2x(2x^2-1)=2x(\sqrt2 x+1)(\sqrt2 x-1)=0 \implies critical~values~x=0,\pm \frac{\sqrt2}{2}\]

OpenStudy (anonymous):

wow okay

OpenStudy (anonymous):

The minimizing critical values are the two nonzero ones.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you understand how this worked out?

OpenStudy (anonymous):

I'm a little bit confused

OpenStudy (anonymous):

About what?

OpenStudy (anonymous):

what you wrote the i see that some numbers didn't show

OpenStudy (anonymous):

Which line?

OpenStudy (anonymous):

the last one where you wrote let f(x) = d(x) and so forth

OpenStudy (anonymous):

All I did was define the function, take the derivative, and set it to zero. On the next line, I factored it to get the roots, which are x=0 and x=plus/minus sqrt2/2.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Keep working hard. Do math every day.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

The final answers will be ordered pairs (-sqrt2/2,1/2) and (sqrt2/2,1/2).